# Fixed-relative-switch strategies for learning based event-triggered control of nonlinear multiagent systems

### v2

@@ Abstract
This paper investigates event-triggered control for consensus tracking in nonlinear semi-strict-feedback multi-agent systems (MASs) with unknown states and subject to disturbances. We begin by employing radial basis function neural networks combined with the backstepping method to approximate the unknown nonlinear dynamics of the MASs, which facilitates the development of state and disturbance observers for estimation. We then propose three adaptive event-triggered control strategies: fixed-threshold, relative-threshold, and switch-threshold, whose controllers and triggering conditions are defined acocrdingly. By applying Lyapunov stability theory, we rigorously prove that all follower agents’ outputs consistently track the reference of their leader, and all error signals remain uniformly bounded. Moreover, our control strategies effectively mitigate the occurrence of Zeno behaviors. To validate our control framework, an illustrative example is provided to demonstrate its desirable performance in consensus tracking and triggering efficiency.
@@
Multiagent system, adaptive control, observer, event-triggered control.
@@ I Introduction
For decades, consensus tracking in nonlinear multi-agent systems (MASs) has been intensively studied in control and learning communities due to its wide applications in fields such as power grids [1][2], intelligent transportation [3][4] and communication systems [5][6]. The primary objective is to ensure that all units within the network operate ”unanimously” along a desired trajectory. However, uncertain system parameters and external perturbations in real engineering scenarios pose significant challenges for the problem and existing methods often fail to work. Consequently, it is critical to develop control strategies that effectively manage unknown states and perturbations.
@@
In the conventional setting of sample-data control system, controllers continuously respond to changes in the system, which often results in unnecessary resource usage, particularly when the network has restricted bandwidth. A pivotal study in 1999 [9] demonstrated the advantages of event-triggered control over periodic impulse control in first-order stochastic systems with multiplicative noise, leading to its widespread adoption in both linear and nonlinear systems. Recent innovative design methodologies, as evidenced by works such as [7]-[16] have significantly advanced the field of event-triggered control.
@@
Event-triggered control was initially applied to nonlinear systems and achieved great success in reducing trigger frequency and enhancing control performance, see, .e.g, [7] and [8]. The framework has then been extended in various directions, where new trigger conditions are introduced. For instance, [7] proposed a simple dynamic triggering condition that alleviates the traditional periodic execution requirements in closed-loop nonlinear systems. And [8] developed a technique that utilizes the current state of the plant to determine triggering time instances, without the need of periodicity. Also event-triggered mechanism has been extensively investigated in the context of MASs more recently. Specifically, [10] introduced a saturated-threshold event-triggered control strategy aimed at achieving robust consensus tracking in continuous-time nonlinear MASs under sensor attacks. And [11] employed a switching-based trigger strategy for multi-agent consensus tracking. Furthermore, [16] analyzed three event-triggered control strategies in nonlinear systems with uncertainty and did not require input-to-state stability (ISS).
Motivated by the above information, we have developed an adaptive consensus tracking framework for high-order nonlinear MASs. Our approach utilizes the backstepping method, filter method, RBF NNs, and observer method. It is specifically designed for scenarios where the states and perturbations are unknown, while only the output signals of the agents are available. The key contributions of this work are outlined below:
  • 1)
@@
    We employ radial basis function neural networks and backstepping method to approximate the unknown transition functions of the MASs under investigation
  • 2)
@@
    Based on the learned model, we develop both state and perturbation observers to monitor unknown states and unmeasured disturbances, which later enhance the robustness of the controller against perturbations.
  • 3)
@@
    We propose three event-triggered control strategies: fixed-threshold, relative-threshold, and switch-threshold, together with a comprehensive analysis of their advantages such as reduced frequency of control updates, enhanced resource utilization and extended controller lifespan.
The organization of this paper is organized in the following way. Section II depicts the system model. The controller and event-triggered strategies are developed in Section III, while Section IV recommended the stability analysis. In Section V, we present our simulation results. Finally, we conclude the paper in Section VI.
II Preliminaries
II-A Problem Formulation
Consider a scenario with $M$ followers, designated as agents $1$ to $M$, and a leader, designated as agent 00, all interacting under a directed communication graph topology. Consider the following class of MASs [19][20]:
@@ II-A Problem Formulation
where $i=1,...,N$, $j=1,...,n$, $k=1,...,n-1$, $\bar{x}_{i,j}=[x_{i,1},...,x_{i,j}]^{T}$ are the state of the $i$th follower, while $u_{i}\in R$ denotes its input of control. The output of the $i$th follower is expressed as $y_{i}\in R$. The function $f_{i,j}(\bar{x}_{i,j})$ represents $C^{1}$ class nonlinear smooth equation vectors. $\xi_{i,j}$ represents unmeasured external perturbations affecting the system. It is noted that The leader’s movement occurs independently, without being influenced by the actions or positions of the followers.
One intention of the works in this paper is to enable each follower’s output signal, ${{y}_{i}}$, to synchronously track the reference signal, ${{y}_{r}}$, through a suitable ETC strategy.
@@ II-B Observer Design
In this paper, to effectively spot the unknown states and perturbations throughout the entire control system, we introduce a series of observers. By precisely analyzing and processing feedback data, these observers can accurately identify and estimate unknown variables within the system, thereby enhancing the overall robustness and stability.
The state observer is defined as
@@
where $\otimes$ represents the Kronecker product, ${{\hat{\bar{x}}}_{i,l}}={{\left[\hat{x}_{i,1}^{T},\ldots,\hat{x}_{i,l}^{T}\right]}^{T}}$ represents the estimated value of the actual state. The vector $Q_{i}=[q_{i,1},...,q_{i,n}]^{T}$ is such that the matrix $P_{i}$ is a strictly Hurwitz. The parameters are defined as follows: $R_{i,1}=\left[\underbrace{0,\ldots,1}_{l},\ldots,0\right]_{n\times 1}^{T}$, $U_{i}=[0,...,0,1]^{T}_{n\times 1}$, $V_{i}=[1,0,...,0]^{T}_{n\times 1}$, $P_{i}=\left[\begin{gathered}-{q_{i,1}}\hfill\\ {\text{ }}\vdots{\text{ }}{I_{(n-1)\times(n-1)}}\hfill\\ -{q_{i,n}}{\text{ 0 }}\ldots{\text{ 0 }}\hfill\\ \end{gathered}\right]$.
Define $\psi_{i}=[(x_{i,1}-\hat{x}_{i,1})^{T},...,(x_{i,n}-\hat{x}_{i,n})^{T}]^{T}$ as the error of state observation, one has that
@@
Define ${{\varphi}_{i}}(t)=\sum\limits_{l=1}^{n}{{{R}_{i,l}}\otimes({{f}_{i,l}}\left({{{\bar{x}}}_{i,l}}\left(t\right)\right)-{{{\hat{f}}}_{i,l}}({{{\hat{\bar{x}}}}_{i,l}}(t)))}$ as the error of function approximation and ${{\varphi}_{i}}(t)={{\left[\varphi_{i,1}^{T}(t),...,\varphi_{i,n}^{T}(t)\right]}^{T}}$, assume that ${{\varphi}_{i}}(t)$ is bounded, this implies that there exists an unspecified parameter $\varphi_{i}^{0}>0$ such that the inequality $\left\|{{\varphi_{i}}(t)}\right\|\leq\varphi_{i}^{0}$ holds true.
By utilizing RBF NNs to approximate the unidentified nonlinear function within the MASs (1), an optimal weight vector $W_{i,l}^{*}$ is defined as:
@@
Consider $\hat{W}_{i,l}^{T}$ as the estimation of $W_{i,l}^{*T}$, the unidentified smooth nonlinear function can be estimated as
${{\hat{f}}_{i,l}}({{\hat{\bar{x}}}_{i,l}})=\hat{W}_{i,l}^{T}{E_{i,l}}\left({{{\hat{\bar{x}}}}_{i,l}}\right)$ (5)
The optimal weight $W_{i,l}^{*}$, with $l=1,\ldots,n$, is designed as follows
@@
The state observers can be revised as follows
@@
with $k=1,...,n-1$.
Define the estimate of external perturbations as $\hat{\varpi}_{i}=[\hat{\varpi}_{i,1}^{T},...,\hat{\varpi}_{i,n}^{T}]^{T}$. and since the external perturbations are unknown, with the aim of achieving consensus control, a novel auxiliary variable is defined as
${{\tau}_{i,l}}={{\varpi}_{i,l}}-{{\kappa}_{i,l}}{{x}_{i,l}}$ (8)
where $\kappa_{i,l}$ is a designed positive parameter.
The perturbation observer is designed as
@@
where ${{x}_{i,n+1}}={{u}_{i}}$ and $\hat{W}_{i,l}$ represents the estimation of $W_{i,l}^{*}$.
II-C Graph Theory
This article examines MASs comprising $\mathcal{M}$ follower agents and one leaders. The interaction among these agents is represented using a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ represents the set of nodes, $\mathcal{V}=\{1,...,N\}$, and $\mathcal{E}$ represents the set of edges, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$. By the way, the set $\mathcal{E}=(\mathcal{V}_{i},\mathcal{V}_{j})\in\mathcal{E}$ means that agent $j$ can gain information from agent $i$. The adjacent matrix $\mathcal{A}=[a_{i,j}]\in R^{N\times N}$, where
@@ II-C Graph Theory
Define the Laplacian matrix $\mathcal{L}$ as $\mathcal{L}=\mathcal{D}-\mathcal{A}$, where $\mathcal{D}=diag({{d_{1}},...,{d_{N}}})$ is defined as the in-degree matrix of graph $\mathcal{G}$, ${d_{i}}=\sum\nolimits_{j=1,j\neq i}^{N}{{a_{ij}}}$. And clarify the diagonal matrix $\mathcal{B}=diag\left({{\mathcal{C}_{1}},...,{\mathcal{C}_{N}}}\right)$, if the leader 0 is capable of transmitting information to agent $i$, then ${\mathcal{C}_{i}}>0$. If not, ${\mathcal{C}_{i}}=0$ holds.
The previous section detailed the system setup process, while Fig. 1 presents the block diagram of the adaptive observer-based control framework with event-triggered strategies.
Figure 1: Block diagram of the adaptive observer-based control with event-triggered strategies.
$Lemma\text{ 1}$[17]: For any specified positive definite matrix ${{H}_{i}}=H_{i}^{T}>0$, with a symmetric positive matrix ${{F}_{i}}$, the equality $B_{i}^{T}{{F}_{i}}+{{F}_{i}}{{B}_{i}}=-2{{H}_{i}}$ holds.
$Lemma\text{ 2}$[18]: The inequality $0\leq\left|\chi\right|-\chi\tanh(\frac{\chi}{{{\chi}_{0}}})\leq 0.2785{{\chi}_{0}}$ always holds with any specified parameter $\chi\in R$ and ${{\chi}_{0}}>0$.
$Assumption\text{ }1$[19] : The unmeasured external perturbations are all bounded, i.e., the inequality $\left\|{{\varpi}_{i,j}}\right\|\leq\varpi_{i,j}^{0}$ holds, where $\varpi_{i,j}^{0}$ defines an positive parameter.
$Assumption\text{ }2$[20] : In this MASs, the leader’s desired reference signal is both measurable and smooth, with ${{y}_{r}}\left(t\right)$, ${{\dot{y}}_{r}}\left(t\right)$ being bounded.
III The Controller Design
Define the graph-based errors $z_{i,k}$ and the boundary layer errors $e_{i,k}$ for the ith follower as
@@ III The Controller Design
$Step\text{ }1$: The differential of $z_{i,1}$ is
${{\dot{z}}_{i,1}}=$ $\left({{d}_{i}}+{{b}_{i}}\right)\left({{{\hat{x}}}_{i,2}}+{{\psi}_{i,2}}\left(t\right)+{{\varpi}_{i,1}\left(t\right)+{{f}_{i,1}}\left({{{\bar{x}}}_{i,1}}\right)}\right)-{{b}_{i}}{{\dot{y}}_{r}}$
$-\sum\limits_{j=1}^{N}{{{a}_{i,j}}\left({{{\hat{x}}}_{j,2}}+{{\psi}_{j,2}}\left(t\right)+{{\varpi}_{j,1}}\left(t\right)+{{f}_{j,1}}\left({{{\bar{x}}}_{j,1}}\right)\right)}$
where $d_{i}=\sum_{j=1}^{N}a_{ij}$.
The Lyapunov function is determined as
@@
where $\psi_{i}=[\psi_{i,1}^{T},...,\psi_{i,n}^{T}]^{T}$, $\eta_{i,1}>0$ and $o_{i}>0$ define designed parameters.
Define an auxiliary function as
@@
${{\bar{Z}}_{i,1}}\left({{T}_{i,1}}\right)$ is estimated by RBF NNs:
${{\bar{Z}}_{i,1}}\left({{T}_{i,1}}\right)=$ $K_{i,1}^{*T}{{E}_{i,1}}\left({{T}_{i,1}}\right)+{{\delta}_{i,1}}\left({{T}_{i,1}}\right)$
$\left|{{\delta}_{i,1}}\left({{T}_{i,1}}\right)\right|\leq$ $\bar{\delta}_{i,1}$ (16)
where $\bar{\delta}_{i,1}$ defines an unknown positive parameter.
Based on the Young’s inequality, the subsequent inequality can be derived:
@@
$+\frac{c_{i,1}^{2}}{2}+\frac{z_{i,1}^{2}}{2}+\frac{\bar{\delta}_{i,1}^{2}}{2}.$ (17)
where ${c}_{i,1}$ is a positive designed parameter and ${{T}_{i,1}}={{[{{y}_{r}},{{\dot{y}}_{r}},\hat{x}_{i,1}^{T},\hat{x}_{j,1}^{T},{{\hat{\varpi}}_{i,1}}^{T},{{\hat{\varpi}}_{j,1}}^{T}]}^{T}}$ denotes the input of RBF NNs.
In accordance with the Young’s inequality and Assumption 1, one has
$-{{\kappa}_{i,1}}{{\tilde{\tau}}_{i,1}}{{\sigma}_{i,1}}\leq$ $\frac{1}{2}\tilde{\tau}_{i,1}^{2}+\frac{\kappa_{i,1}^{2}\sigma_{0,1}^{2}}{2}$ (18)
$-{{\kappa}_{i,1}}{{\tilde{\tau}}_{i,1}}{{\psi}_{i,2}}\leq$ $\frac{1}{2}\tilde{\tau}_{i,1}^{2}+{{{\kappa_{i,1}^{2}\left\|{{\psi}_{i}}\right\|}^{2}}}$ (19)
$-\kappa_{i,1}^{2}{{\tilde{\tau}}_{i,1}}{{\psi}_{i,1}}\leq$ $\frac{1}{2}\tilde{\tau}_{i,1}^{2}+{{{\kappa_{i,1}^{4}\left\|{{\psi}_{i}}\right\|}^{2}}}$ (20)
@@
The design of the virtual control law and the adaptive law is formulated as
${{\alpha}_{i,2}}=$ $\frac{1}{{{d}_{i}}+{{b}_{i}}}({{r}_{i,1}}{{z}_{i,1}}-\frac{{{z}_{i,1}}}{2}-\frac{{{{\hat{\Theta}}}_{i}}}{2c_{i,1}^{2}}z_{i,1}E_{i,1}^{T}\left({{T}_{i,1}}\right){{E}_{i,1}}\left({{T}_{i,1}}\right))$ (23)
@@
where ${{h}_{i,1}}$ and ${{r}_{i,1}}$ are designed parameters.
Then, considering Lemma 1 and the above formulas, one has
@@
where ${{\iota}_{i,1}}=\frac{c_{i,1}^{2}}{2}+\frac{\bar{\delta}_{i,1}^{2}}{2}+\frac{\kappa_{i,1}^{2}\sigma_{0,1}^{2}}{2}+\frac{1}{2}{{\left\|\varphi_{i}^{0}\right\|}^{2}}{{\left\|{{F}_{i}}\otimes{{I}_{m}}\right\|}^{2}}+\frac{1}{2}{{\left\|\tilde{\varpi}_{i}^{0}\right\|}^{2}}{{\left\|{{F}_{i}}\otimes{{I}_{m}}\right\|}^{2}}$.
Then, process ${{\alpha}_{i,2}}$ through the first-order low-pass filter, resulting in ${{\bar{\alpha}}_{i,2}}$ as follows:
@@
where $m_{i,2}$ is a small positive designed parameter.
$Step\text{ }k$: Differentiating ${{z}_{i,k}}$, one has
${{\dot{z}}_{i,k}}=$ ${{z}_{i,k+1}}+{{s}_{i,k+1}}+\hat{W}_{i,k}^{T}{{E}_{i,k}}\left({{{\hat{\bar{x}}}}_{i,k}}\right)+{{q}_{i,k}}{{\psi}_{i,1}}$
$+{{\hat{\varpi}}_{i,k}}(t)+{{\alpha}_{i,k+1}}-{{\dot{\bar{\alpha}}}_{i,k}}$ (27)
where $k=2,...,n-1$.
Choose the Lyapunov function as
@@
where ${{\eta}_{i,k}}$ is a positive parameter.
By utilizing RBF NNs, one has
${{\bar{Z}}_{i,k}}\left({{T}_{i,k}}\right)=$ $K_{i,k}^{*T}{{E}_{i,k}}\left({{T}_{i,k}}\right)+{{\delta}_{i,k}}\left({{T}_{i,k}}\right)$
$\left|{{\delta}_{i,k}}\left({{T}_{i,k}}\right)\right|\leq$ $\bar{\delta}_{i,k}$ (29)
where $\bar{\delta}_{i,k}$ denotes an unknown positive parameter.
The design of the virtual control law and the adaptive law is formulated as follows
@@
where ${{T}_{i,k}}={{[{{y}_{r}},{{\hat{\Theta}}_{i}},\hat{\bar{x}}_{i,k}^{T},\hat{\bar{x}}_{j,k}^{T},\hat{\varpi}_{i,k}^{T},\hat{\varpi}_{j,k}^{T}]}^{T}}$, ${{r}_{i,k}}$ and ${{h}_{i,k}}$ are designed parameters.
In terms of the Young’s inequality, one has
@@
where ${{\iota}_{i,k}}=\frac{c_{i,k}^{2}}{2}+\frac{\bar{\delta}_{i,k}^{2}}{2}+\frac{\kappa_{i,k}^{2}\sigma_{0,k}^{2}}{2}$, and $c_{i,k}$ denotes a positive parameter.
In line with Step 1, the first-order low-pass filter is implemented as follows:
@@
${{m}_{i,k+1}}{{\dot{\bar{\alpha}}}_{i,k+1}}+{{\bar{\alpha}}_{i,k+1}}={{\alpha}_{i,k+1}}.$ (33)

### v3

# Fixed-Relative-Switched Threshold Strategies for Consensus Tracking Control of Nonlinear Multiagent Systems

@@ Abstract
This paper investigates event-triggered consensus tracking in nonlinear semi-strict-feedback multi-agent systems involving one leader and multiple followers. We first employ radial basis function neural networks and backstepping techniques to approximate the unknown nonlinear dynamics, facilitating the design of dual observers to measure the unknown states and disturbances. Then three adaptive event-triggered control schemes are proposed: fixed-threshold, relative-threshold, and switched-threshold configurations, each featuring specialized controller architectures and triggering mechanisms. Through Lyapunov stability analysis, we establish that the follower agents can asymptotically track the reference trajectory of the leader, meanwhile all error signals remain uniform bounded. Our proposed control strategies effectively prevent Zeno behaviors through stringent exclusion criteria. Finally, an illustrative example is presented, demonstrating the competitive performance of our control framework in achieving consensus tracking and optimizing triggering efficiency.
@@
Multiagent systems, consensus tracking, observer design, event-triggered control, learning-based control
@@ I Introduction
For decades, consensus tracking in nonlinear multi-agent systems (MASs) has been intensively studied in control and learning communities due to its wide applications such as power grids [1, 2], intelligent transportation [3, 4, 5] and communication systems [6, 7]. The primary objective is to ensure that all units in the network operate “unanimously” along a desired trajectory and achieve consensus [8]. However, the presence of uncertain system parameters and external perturbations in real engineering scenarios poses considerable challenges for consensus tracking, often rendering existing methods fail to work. Consequently, it is critical to develop control methods that effectively address the issues.
@@
In the conventional settings of the sample-data control system, controllers continuously respond to changes in the system, which often results in unnecessary resource usage, particularly when the network has restricted bandwidth. A seminal work [11] demonstrated the advantages of event-triggered control over periodic impulse control in first-order stochastic systems with multiplicative noise, leading to its widespread adoption in both linear and nonlinear systems. Recent advancements in controller design such as [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19] have further propelled the field of event-triggered control.
@@
The application of event-triggered control to nonlinear systems has achieved notable success in reducing trigger frequency and enhancing control performance [9, 10]. This framework has since been extended by introducing new triggering conditions. For instance, [9] proposed a dynamic triggering condition that alleviates the traditional periodic execution requirements in closed-loop nonlinear systems. The work in [10] utilized the current state of the plant to determine triggering time instances without the requirement of periodicity. More recently, event-triggered mechanisms have been extensively explored within the context of MASs. Notably, [12] introduced a saturated-threshold event-triggered control strategy which aims to achieve robust consensus tracking in continuous-time nonlinear MASs amidst sensor attacks, and [13] employed a switching-based trigger strategy for consensus tracking. Additionally, [18] analyzed three event-triggered control strategies in nonlinear uncertain systems, which did not require input-to-state stability (ISS).
In light of these results, we propose an adaptive consensus tracking framework for high-order nonlinear MASs, which integrates the backstepping method, filtering techniques, radial basis function neural networks (RBF NNs), and observer design. Our key contributions are summarized as follows:
@@
  • We employ RBF NNs with backstepping method to approximate the unknown transition functions of the MASs;
@@
  • We develop observers to monitor both unknown states and unmeasured disturbances, which later enhance the robustness of the controller against perturbations;
@@
  • We propose three event-triggered control strategies: fixed-threshold, relative-threshold, and switched-threshold, then highlight their performance in terms of reduced control update frequency, improved resource utilization, and extended controller lifespan in the empirical study.
The remainder of the work is structured as follows. Section II reviews preliminary knowledge of MAS, observer design and graph theory, then formulates the key problem of the work. Section III develops the fixed-relative-switched threshold control strategies and analyzes the stability of the strategies. Section IV presents simulation results to validate the performance of our method. Finally, Section V concludes the paper and lists several future research directions.
II Preliminaries and Problem Formulation
II-A System Model
Consider a team of agents with $N$ followers indexed from $1$ to $N$ and a leader indexed 00, they communicate with each other in a directed graph and form the MAS model [22] [23]:
@@ II-A System Model
where $i=1,...,N$, $j=1,...,n$, $k=1,...,n-1$, $\bar{x}_{i,j}=[x_{i,1},...,x_{i,j}]^{T}$ are the state of the $i$th follower, while $u_{i}\in R$ denotes its input of control. The output of the $i$th follower is expressed as $y_{i}\in R$. The function $f_{i,j}(\bar{x}_{i,j})$ represents $C^{1}$ class nonlinear smooth equation vectors. $\xi_{i,j}$ represents unmeasured external perturbations affecting the system. Notably, the leader’s movement occurs independently without being influenced by the actions or positions of the followers.
@@ II-B Observer Design
To effectively locate the unknown states and perturbations throughout the entire control system, we introduce a series of observers. By precisely analyzing and processing feedback data, these observers accurately identify and estimate unknown variables within the system, enhancing the overall robustness and stability. The state observer is defined as
@@
where $\otimes$ represents the Kronecker product, and ${{\hat{\bar{x}}}_{i,l}}={{[\hat{x}_{i,1}^{T},\ldots,\hat{x}_{i,l}^{T}]}^{T}}$ represents the estimated value of the actual state with $l=1,...,n$. The vector $Q_{i}=[q_{i,1},...,q_{i,n}]^{T}$ is such that the matrix $P_{i}$ is a strictly Hurwitz. The parameters are defined as follows: $R_{i,l}=[0,\ldots,1,\ldots,0]_{n\times 1}^{T}$ where the $l$-th element is $1$, $U_{i}=[0,...,0,1]^{T}_{n\times 1}$, $V_{i}=[1,0,...,0]^{T}_{n\times 1}$ and $P_{i}=[-Q_{i},R_{i,1},\ldots,R_{i,n-1}]$.
We denote by $\psi_{i}=[(x_{i,1}-\hat{x}_{i,1})^{T},...,(x_{i,n}-\hat{x}_{i,n})^{T}]^{T}$ the error of state observation and have that
@@
We also denote the error of function approximation by ${{\varphi}_{i}}(t)=\sum\limits_{l=1}^{n}{{{R}_{i,l}}\otimes({{f}_{i,l}}\left({{{\bar{x}}}_{i,l}}\left(t\right)\right)-{{{\hat{f}}}_{i,l}}({{{\hat{\bar{x}}}}_{i,l}}(t)))}$ and write ${{\varphi}_{i}}(t)={{\left[\varphi_{i,1}^{T}(t),...,\varphi_{i,n}^{T}(t)\right]}^{T}}$. Suppose that ${{\varphi}_{i}}(t)$ is bounded, thus there exists an unspecified parameter $\varphi_{i}^{0}>0$ such that the inequality $\left\|{{\varphi_{i}}(t)}\right\|\leq\varphi_{i}^{0}$ holds. Using RBF NNs to approximate the unidentified nonlinear function of the MASs (II-A), an optimal weight vector $W_{i,l}^{*}$ is derived as:
@@
Let $\hat{W}_{i,l}^{T}$ be the estimation of $W_{i,l}^{*T}$, then the estimated smooth nonlinear function is ${{\hat{f}}_{i,l}}({{\hat{\bar{x}}}_{i,l}})=\hat{W}_{i,l}^{T}{E_{i,l}}\left({{{\hat{\bar{x}}}}_{i,l}}\right)$. The optimal weight $W_{i,l}^{*}$ for $l=1,\ldots,n$ is designed as:
@@
The state observers is then revised as: $\forall 1\leq k<n$
@@
Next, we define the estimate of unknown external perturbations as $\hat{\varpi}_{i}=[\hat{\varpi}_{i,1}^{T},...,\hat{\varpi}_{i,n}^{T}]^{T}$. In order to achieve consensus control, we introduce an auxiliary variable ${\tau}_{i,l}={\varpi}_{i,l}-{\kappa}_{i,l}{x}_{i,l}$ where $\kappa_{i,l}$ is a positive parameter.
Given the above concepts, the perturbation observer is:
@@
where ${{x}_{i,n+1}}={{u}_{i}}$ and $\hat{W}_{i,l}$ is the estimation of $W_{i,l}^{*}$.
II-C Directed Graph
Given MASs with $N$ followers and one leader, the interaction between agents is modeled by a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{1,...,N\}$ represents the set of nodes, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ represents the set of edges, and $(\mathcal{V}_{i},\mathcal{V}_{j})\in\mathcal{E}$ indicates that agent $j$ can gain information from agent $i$. The adjacent matrix $\mathcal{A}=[a_{i,j}]\in R^{N\times N}$ of $\mathcal{G}$ is defined as:
@@ II-C Directed Graph
We introduce the Laplacian matrix $\mathcal{L}$ as $\mathcal{L}=\mathcal{D}-\mathcal{A}$, where $\mathcal{D}=diag({{d_{1}},...,{d_{N}}})$ is the in-degree matrix of graph $\mathcal{G}$, ${d_{i}}=\sum\nolimits_{j=1,j\neq i}^{N}{{a_{ij}}}$. And clarify the diagonal matrix $\mathcal{B}=diag\left({{\mathcal{C}_{1}},...,{\mathcal{C}_{N}}}\right)$, if the leader 0 is capable of transmitting information to agent $i$, then ${\mathcal{C}_{i}}>0$. If not, ${\mathcal{C}_{i}}=0$ holds.
II-D Problem Formulation
Figure 1: Block diagram of our proposed control framework.
Fig. 1 illustrates our framework of adaptive event-triggered consensus tracking control. The following lemmas and assumptions are necessary to facilitate the later discussions.
Lemma 1 (from [20]).
For any positive definite matrix ${{H}_{i}}=H_{i}^{T}>0$, with a symmetric positive matrix ${{F}_{i}}$ and a strict Hurwitz matrix $P_{i}$, it satisfies $P_{i}^{T}{{F}_{i}}+{{F}_{i}}{{P}_{i}}=-2{{H}_{i}}$.
Lemma 2 (from [21]).
Inequality $0\leq\left|\chi\right|-\chi\tanh(\frac{\chi}{{{\chi}_{0}}})\leq 0.2785{{\chi}_{0}}$ holds for any parameter $\chi\in R$ and ${{\chi}_{0}}>0$.
Assumption 1 (from [15, 22, 25]).
The unmeasured external perturbations are all bounded, i.e., inequality $\left\|{{\varpi}_{i,j}}\right\|\leq\varpi_{i,j}^{0}$ holds, where $\varpi_{i,j}^{0}$ is a positive parameter.
Assumption 2 (from [23]).
The desired reference signal of the leader in the MASs is both measurable and smooth, also ${{y}_{r}}\left(t\right)$ and its first order derivative ${{\dot{y}}_{r}}\left(t\right)$ are bounded.
Problem 1 (Consensus tracking of MASs).
Given a class of nonlinear MASs following equation (II-A), suppose that the communication of agents is modeled by a directed graph $\mathcal{G}$, each follower is subject to unknown external perturbations $\xi$ and required to track the leader’s reference signal $y_{r}$, then our goals are two-fold: (i) design an adaptive control law to ensure that the tracking errors are uniformly bounded; (ii) evaluate the performance of different event-triggered threshold control strategies in terms of control update frequency, resources conservation and controller’s lifespan.
III Event-Triggered Controller Design
This section is initiated by the general procedure of controller design for Problem 1. First, the graph-based errors $z_{i,k}$ and boundary layer errors $e_{i,k}$ for the $i$-th follower are:
@@ III Event-Triggered Controller Design
$Step\text{ }1$: The Lyapunov function is determined as
@@
where $\psi_{i}=[\psi_{i,1}^{T},...,\psi_{i,n}^{T}]^{T}$, $\eta_{i,1}>0$ and $o_{i}>0$ define designed parameters. Then, define an auxiliary function as
@@
${{\bar{Z}}_{i,1}}\left({{T}_{i,1}}\right)$ is estimated by RBF NNs: ${{\bar{Z}}_{i,1}}\left({{T}_{i,1}}\right)=K_{i,1}^{*T}{{E}_{i,1}}\left({{T}_{i,1}}\right)+{{\delta}_{i,1}}\left({{T}_{i,1}}\right)$, where $\bar{\delta}_{i,1}$ defines an unknown positive parameter, ${{T}_{i,1}}={{[{{y}_{r}},{{\dot{y}}_{r}},\hat{x}_{i,1}^{T},\hat{x}_{j,1}^{T},{{\hat{\varpi}}_{i,1}}^{T},{{\hat{\varpi}}_{j,1}}^{T}]}^{T}}$ denotes the input of RBF NNs and $K_{i,1}^{*T}$ is the optimal weight, just like the $W_{i,1}^{*T}$ in (4)(5), with $\left|{{\delta}_{i,1}}\left({{T}_{i,1}}\right)\right|\leq\bar{\delta}_{i,1}$.
Based on the Young’s inequality, with a positive designed parameter ${c}_{i,1}$, the subsequent inequality can be derived:
@@
$+\frac{c_{i,1}^{2}}{2}+\frac{z_{i,1}^{2}}{2}+\frac{\bar{\delta}_{i,1}^{2}}{2}$ (12)
By Young’s inequality and Assumption 1, one has
@@
The virtual and adaptive control laws are formulated as
${{\alpha}_{i,2}}=$ $\frac{1}{{{d}_{i}}+{{b}_{i}}}({{r}_{i,1}}{{z}_{i,1}}-\frac{{{z}_{i,1}}}{2}$
$-\frac{{{{\hat{\Theta}}}_{i}}}{2c_{i,1}^{2}}z_{i,1}E_{i,1}^{T}\left({{T}_{i,1}}\right){{E}_{i,1}}\left({{T}_{i,1}}\right))$ (15)
@@
where ${{h}_{i,1}}$ and ${{r}_{i,1}}$ are designed parameters. Then, considering Lemma 1 and the above formulas, one has
@@
where ${{\iota}_{i,1}}=\frac{c_{i,1}^{2}}{2}+\frac{\bar{\delta}_{i,1}^{2}}{2}+\frac{\kappa_{i,1}^{2}\sigma_{0,1}^{2}}{2}+\frac{1}{2}{{\left\|\varphi_{i}^{0}\right\|}^{2}}{{\left\|{{F}_{i}}\otimes{{I}_{m}}\right\|}^{2}}+\frac{1}{2}{{\left\|\tilde{\varpi}_{i}^{0}\right\|}^{2}}{{\left\|{{F}_{i}}\otimes{{I}_{m}}\right\|}^{2}}$. Then, with a small positive designed parameter $m_{i,2}$, process ${{\alpha}_{i,2}}$ through the first-order low-pass filter, resulting in ${{\bar{\alpha}}_{i,2}}$ as follows:
@@
$Step\text{ }k$: Choose the Lyapunov function as
@@
where ${{\eta}_{i,k}}>0$ is a parameter. By utilizing RBF NNs, one has ${{\bar{Z}}_{i,k}}\left({{T}_{i,k}}\right)=K_{i,k}^{*T}{{E}_{i,k}}\left({{T}_{i,k}}\right)+{{\delta}_{i,k}}\left({{T}_{i,k}}\right)$, where $\bar{\delta}_{i,k}>0$ is an unknown parameter, with $|{\delta}_{i,k}(T_{i,k})|\leq\bar{\delta}_{i,k}$. Then we design the virtual control law and adaptive law as:
@@
where ${{T}_{i,k}}={{[{{y}_{r}},{{\hat{\Theta}}_{i}},\hat{\bar{x}}_{i,k}^{T},\hat{\bar{x}}_{j,k}^{T},\hat{\varpi}_{i,k}^{T},\hat{\varpi}_{j,k}^{T}]}^{T}}$, ${{r}_{i,k}}$ and ${{h}_{i,k}}$ are designed parameters. By the Young’s inequality, one has
@@
where ${{\iota}_{i,k}}=\frac{c_{i,k}^{2}}{2}+\frac{\bar{\delta}_{i,k}^{2}}{2}+\frac{\kappa_{i,k}^{2}\sigma_{0,k}^{2}}{2}$ and $c_{i,k}0>0$. In line with Step 1, the first-order low-pass filter is implemented as:
@@
${{m}_{i,k+1}}{{\dot{\bar{\alpha}}}_{i,k+1}}+{{\bar{\alpha}}_{i,k+1}}={{\alpha}_{i,k+1}}$ (23)